### v1

# Opinion Dynamics with Multiple Adversaries

Abstract.
Opinion dynamics model how the publicly expressed opinions of users in a social network coevolve according to their neighbors as well as their own intrinsic opinion. Motivated by the real-world manipulation of social networks during the 2016 US elections and the 2019 Hong Kong protests, a growing body of work models the effects of a strategic actor who interferes with the network to induce disagreement or polarization. We lift the assumption of a single strategic actor by introducing a model in which any subset of network users can manipulate network outcomes. They do so by acting according to a fictitious intrinsic opinion. Strategic actors can have conflicting goals, and push competing narratives. We characterize the Nash Equilibrium of the resulting meta-game played by the strategic actors. Experiments on real-world social network datasets from Twitter, Reddit, and Political Blogs show that strategic agents can significantly increase polarization and disagreement, as well as increase the “cost” of the equilibrium. To this end, we give worst-case upper bounds on the Price of Misreporting (analogous to the Price of Anarchy). Finally, we give efficient learning algorithms for the platform to (i) detect whether strategic manipulation has occurred, and (ii) learn who the strategic actors are. Our algorithms are accurate on the same real-world datasets, suggesting how platforms can take steps to mitigate the effects of strategic behavior.
conference: ; ;
1. Introduction
Over the past decade, social media has experienced rapid growth in both usage and significance. Online social networks, which allow users to share updates about their lives and opinions with a broad audience instantaneously, are now utilized by billions of people globally. These platforms serve various purposes, such as being informed about politics, news, health-related updates, products, and many more (Backstrom et al., 2012; Young, 2006; Banerjee et al., 2013; Shearer and Mitchell, 2021).
Unfortunately, networks can induce polarization, with the network connections serving as a pathway for social discord to increase (Musco et al., 2018a; Chen and Rácz, 2021a; Wang and Kleinberg, 2024; Gaitonde et al., 2020). This is a well-studied sociological phenomenon called the filter-bubble theory (Pariser, 2011). The filter-bubble theory argues that personalized algorithms used by online platforms, such as search engines and social media, selectively display content that aligns with a user’s past behaviors, preferences, and beliefs. This customization creates an “invisible algorithmic editing” of the web, isolating individuals within their own ideological bubbles where they encounter only information that reinforces their existing views. As a result, users are less likely to be exposed to diverse perspectives, potentially narrowing their worldview and fostering polarization. Pariser (2011) warns that such bubbles undermine democratic discourse by limiting opportunities for individuals to engage with challenging or unfamiliar ideas.
Additionally, social networks can be manipulated by malicious entities in order to create discord and cause disagreement. For instance, the 2017 indictment of the Russian Internet Research Agency (IRA) by the U.S. Department of Justice Special Counsel’s Office alleged that the IRA leveraged multiple social media accounts and targeted advertising to achieve “a strategic goal to sow discord in the U.S. political system, including the 2016 U.S. presidential election” (Mueller, 2018). In 2019, Twitter, Inc. (2019) disclosed that at least 936 accounts attempted to induce discord in Hong Kong, to e.g. hinder protesters’ ability to organize effectively during the independence movement. As social media continues to proliferate, it is likely that these types of external interferences will become increasingly common. Additionally, networks of Facebook pages have targeted Americans with sports betting scams, amplifying their reach by disseminating provocative conspiracy theories about political figures and natural disasters (Bjork-James and Donovan, 2024). These schemes leverage the economics of the internet, where engagement with inflammatory content is monetized, and social media algorithms inadvertently amplify such content, enabling bad actors to exploit audiences for profit.
To model the opinions’ evolution, computer scientists, sociologists, and statisticians have relied on the framework of opinion dynamics where the users’ opinions coevolve according to a weighted network $G=(V,E,w)$, and each user updates their opinion as a combination of their own intrinsic opinion as well as the opinions of their neighbors (Friedkin and Johnsen, 1990). This model of opinion exchange has the advantage of taking into account both network interactions and their own intrinsic opinion. So far, all of the existing works consider a single actor who has the ability to act on the network to induce disagreement or polarization (Musco et al., 2018a; Chen and Rácz, 2021a; Wang and Kleinberg, 2024; Ristache et al., 2024; Gaitonde et al., 2020; Rácz and Rigobon, 2023; Chitra and Musco, 2020).
Figure 1. Visualization of the strategic equilibrium ($\bm{z}^{\prime}$) on the Karate Club Graph for two different choices of $S$. The truthful intrinsic opinions have been taken to be $\bm{s}=\bm{u}_{2}$ where $\bm{u}_{2}$ is the Fiedler eigenvector of $G$. The white nodes correspond to the nodes in $S$. For the other nodes, the nodes colored in blue (resp. red) correspond to nodes whose public opinion $\bm{z}^{\prime}_{i}$ increased (resp. decreased), i.e., $(\bm{z}^{\prime}_{i}-\bm{z}_{i})/\bm{z}_{i}\geq 0$ (resp. $(\bm{z}^{\prime}_{i}-\bm{z}_{i})/\bm{z}_{i}<0$) after $\bm{s}^{\prime}$ was chosen.
In this work, we lift the assumption of requiring a single actor (such as the platform) to act as an adversary to induce polarization or disagreement and consider the case of several decentralized actors. It is known that empirically, a very small percentage ($25\%$) of the users in a network need to disagree to sway consensus (Centola et al., 2018). Moreover, real-world social networks involve multiple malicious actors, who use different levels of manipulation and hate speech based on their individual goals (Bjork-James and Donovan, 2024). In this paper, we attempt to provide a theoretical basis for this phenomenon: specifically, in our setting, we assume that there is a set $S\subseteq V$ of strategic agents whose goal is to report false intrinsic opinions ($\bm{s}^{\prime}$) that are different from their true intrinsic opinions ($\bm{s}\neq\bm{s}^{\prime}$). Their goal is to influence others while not deviating much from their neighbors; namely, they want to reach an equilibrium where their neighbors agree with them.
For instance, assume a social network where a set of $S$ of political actors want the network to believe that their stance on a topic (e.g., abortion, elections, drug legalization, etc.) is the best. They achieve this by adversarially reporting different intrinsic opinions. This ensures that their influence is both persuasive and credible within the local network context. Such adversarial behavior can result in significantly different (cf. Figure 1) and highly polarized equilibria, where the strategic agents’ opinions appear dominant despite not reflecting the actual intrinsic views of the majority.
Our work investigates the conditions under which these strategic manipulations are successful, the extent of their impact on network-wide opinion dynamics, and how platforms can learn from observing these manipulated equilibria to mitigate such impacts.
1.1. Our Contributions
In this paper, we ask the following research question (RQ):
(RQ) What if a set of strategic actors with possibly conflicting goals tries to manipulate the consensus by strategically reporting beliefs different than their true beliefs?
We rely on the Friedkin-Johnsen (FJ) model (Friedkin and Johnsen, 1990), where the opinions of agents coevolve via the help of a weighted undirected network $G=(V=[n],E,w)$ with non-negative weights. According to the FJ model, the agents possess intrinsic opinions $s$ and express opinions $z$, which they update via the following rule for each agent $i$:
(1) $\bm{z}_{i}(t+1)=\frac{\alpha_{i}\bm{s}_{i}+(1-\alpha_{i})\sum_{i\sim j}w_{ij}\bm{z}_{j}(t)}{1+\sum_{i\sim j}w_{ij}}.$
where $\alpha_{i}\in(0,1)$ is $i$’s susceptibility to persuation (Abebe et al., 2018). We additionally define $\tilde{\alpha}_{i}=\alpha_{i}/(1-\alpha_{i})$ to be the normalized susceptibility parameter corresponding to $i$. This update rule corresponds to the best-response dynamics arising from minimizing the quadratic cost function for each $i$ (Bindel et al., 2011; Abebe et al., 2018):
(2) $c_{i}(\bm{z}_{i},\bm{z}_{-i})=(1-\alpha_{i})\sum_{i\sim j}w_{ij}(\bm{z}_{i}-\bm{z}_{j})^{2}+\alpha_{i}(\bm{z}_{i}-\bm{s}_{i})^{2}.$
The Pure Strategy Nash Equilibrium (PSNE) can be written as $z=((I-A)L+A)^{-1}A\bm{s}=B\bm{s}$ where $L$ is the Laplacian of graph $G$, $A=\mathrm{diag}(\alpha_{1},\dots,\alpha_{n})$ is the diagonal matrix of susceptibilities. When an external single actor aims to induce disagreement or polarization – see, e.g., Gaitonde et al. (2020); Racz and Rigobon (2022); Musco et al. (2018a) – the adversary is coined with optimizing the objective function
$\sum_{i\in[n]}c_{i}(\bm{z}_{i},\bm{z}_{-i})=\bm{s}^{T}((I-A)L+A)^{-1}Af(L)((I-A)L+A)^{-1}A\bm{s},$
where $f(L)$ is a function of the Laplacian of $G$, either with optimizing towards $\bm{s}$ (Gaitonde et al., 2020), or the graph itself (Musco et al., 2018a; Racz and Rigobon, 2022).
Usually, as we also discussed earlier, many adverse actions on social networks come from several independent strategic adversaries who try to manipulate the network by infiltrating intrinsic opinions $\bm{s}_{i}^{\prime}$, which are different from their true stances $\bm{s}_{i}$ but are simultaneously close to $\bm{s}_{i}$. Unlike previous works, these “adversaries” can have conflicting goals.
Concretely, the true opinions of the agents are $\bm{s}_{1},...,\bm{s}_{n}\in\mathbb{R}$, and there is a set $S$ of deviating agents who report $\{\bm{s}_{i}^{\prime}\}_{i\in S}$. The goal of the strategic agents is to minimize the cost function of Equation 2 at consensus $z^{\prime}=((I-A)L+A)^{-1}A\bm{s}^{\prime}$ where $\bm{s}^{\prime}$ is the vector which has entries $\bm{s}_{i}$ for all $i\notin S$ and $\bm{s}_{i}^{\prime}$ for all $i\in S$. The local optimization of agent $i$ becomes:
(3) $\min_{\bm{s}_{i}^{\prime}\in\mathbb{R}}c_{i}\left(\bm{z}^{\prime}=((I-A)L+A)^{-1}A\bm{s}^{\prime}\right).$
Our contributions are as follows.
Characterizing Nash Equilibria with Multiple Adversaries.
We give the Nash equilibrium of the game defined by Equation 3, and show that all Nash-optimal strategies are pure. The Pure Strategy Nash Equilibrium (PSNE) that is given by solving a constrained linear system. Given the PSNE of the game, we characterize the actors who can have the most influence in strategically manipulating the network.
Real-World Experiments to Understand Properites of Equilibria.
We apply our framework to real-world social network data from Twitter and Reddit (Chitra and Musco, 2020), and data from the Political Blogs (Polblogs) dataset (Adamic and Glance, 2005). We find that the influence of strategic agents can be rather significant as they can significantly increase polarization and disagreement, as well as increase the overall “cost” of the consensus.
Analysis of Equilibrium Outcomes Under Different Sets of Strategic Actors.
Various metrics for network polarization and disagreement are sensitive to the choice of who acts strategically, in nontrivial ways. For example, adding more strategic agents can sometimes decrease the Disagreement Ratio at equilibrium (Figure 5), due to counterbalancing effects. To address the effects of manipulation, we give worst-case upper bounds on the Price of Misreporting (PoM), which is analogous to well-studied Price of Anarchy bounds (see, for example, Bhawalkar et al. (2013); Roughgarden and Schoppmann (2011)), and suggest ways that the platform can be used to mitigate the effect of strategic behavior on their network.
Learning Algorithms for the Platform.
We give an efficient algorithm for the platform to detect if manipulation has occurred (Algorithm 1), based on a hypothesis test with the publicly reported opinions $\bm{z}^{\prime}$. Next, we give an algorithm to infer who manipulated the network (the set of strategic agents $S$) from $\bm{z}^{\prime}$, as long as the size of $S$ is sufficiently small. Our algorithm is inspired by the robust regression algorithm of Bhatia et al. (2015), and is practical for real-world networks. It (i) requires the platform to have access to node embeddings $X$ which have been shown computable even in billion-scale networks such as Twitter (El-Kishky et al., 2022), and (ii) can be computed in time $(n+m)^{O(1)}$, where $n$ is the number of nodes and $m$ is the number of edges of the network. Our algorithms have high accuracy on real-world datasets from Twitter, Reddit, and Polblogs.
1.2. Preliminaries and Notations
The set $[n]$ denotes $\{1,\dots,n\}$. $\|\bm{x}\|_{p}$ denotes the $\ell_{p}$-norm of vector $\bm{x}$, whose $i$-th entry is denoted by $\bm{x}_{i}$. $X\succeq 0$ denotes that the matrix $X$ is positive semi-definite and $\|X\|_{2}$ corresponds to the spectral norm of $X$. The Laplacian of the graph $G$ is denoted by $L=D-W$ where $W$ is the weight matrix of the graph, which has entries $w_{ij}\geq 0$, and $D$ is the diagonal degree matrix with diagonal entries $D_{ii}=\sum_{i\sim j}w_{ij}$. The Laplacian has eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$. For any undirected and connected graph $G$, $L$ is symmetric and PSD, so we can write the eigendecomposition of $L$ as:
(4) $L=\sum_{i\in[n]}\lambda_{i}\bm{u}_{i}\bm{u}_{i}^{T}\succeq 0,$
where $\bm{u}_{1},\dots,\bm{u}_{n}$ are orthonormal eigenvectors. Moreover, $\bm{u}_{1}=(1/\sqrt{n})\mathbf{1}$, where $\mathbf{1}$ is the column vector of all 1s. $U$ denotes the matrix which has the eigenvectors of $L$ as columns; i.e., such that $L=U^{T}\Lambda U$ where $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ is the diagonal matrix of $L$’s eigenvalues. $L_{i}$ denotes the $i$-restricted Laplacian which corresponds to the Laplacian of the graph with all edges that are non-adjacent to $i$ being removed, and, similarly, $L_{\{u,v\}}$ corresponds to the Laplacian of an edge $\{u,v\}$. Note that $L_{i}=\sum_{i\sim j}L_{\{i,j\}}$. For a function $f(L)$ of the Laplacian we write $f(L)=U^{T}f(\Lambda)U$ where $f(\Lambda)=\mathrm{diag}(f(\lambda_{1}),\dots,f(\lambda_{n}))$. For brevity, regarding the equilibrium $z$ of the FJ model, we write $B=((I-A)L+A)^{-1}A$, such that $\bm{z}=B\bm{s}$ and $\bm{z}^{\prime}=B\bm{s}^{\prime}$. Finally $\bm{e}_{1},\dots,\bm{e}_{n}\in\mathbb{R}^{n}$ denote the canonical basis.
We define the total cost of an equilibrium $\bm{z}$ to be
(5) $C(\bm{z})=\sum_{i\in[n]}c_{i}(\bm{z}).$
We define the platform-wide metrics to be
(6) Polarization $\textrm{Polarization Ratio}\qquad\mathcal{P}(\bm{z})$ $=\sum_{i\in[n]}(\bm{z}_{i}-\bar{z})^{2},\text{ where }\bar{z}=\frac{1}{n}\sum_{i\in[n]}\bm{z}_{i},$
(7) Disagreement $\textrm{Disagreement Ratio}\qquad\mathcal{D}(\bm{z})$ $=\sum_{i,j\in[n]}w_{ij}(\bm{z}_{i}-\bm{z}_{j})^{2}=\bm{z}^{T}L\bm{z}.$
Finally, we define the “Price of Misreporting” (PoM), which is analogous to the Price of Anarchy (Roughgarden, 2005). The PoM is the ratio of the cost $C(\bm{z}^{\prime})$ when the agents are deviating, and the cost $C(\bm{z})$ when the agents are reporting truthfully, i.e.,
(8) $\mathsf{PoM}:=\frac{C(\bm{z}^{\prime})}{C(\bm{z})}.$
Unlike the Price of Anarchy (PoA), the equilibrium $\bm{z}$ in the denominator of Eq. (8) is the Nash equilibrium for the Friedkin-Johnson dynamics without manipulation. In the PoA, the denominator would be $C(\bm{z}^{*})$, where $\bm{z}^{*}$ is a socially optimal equilibrium (Bindel et al., 2011). Since we study strategic manipulations as a meta-game with respect to the base game of FJ dynamics, it is more relevant for us to compare $\bm{z}^{\prime}$ with $\bm{z}$ than with $\bm{z}^{*}$. Note that $C(\bm{z}^{*})\leq C(\bm{z})\leq C(\bm{z}^{\prime})$, so $\mathsf{PoA}\geq\mathsf{PoM}\geq 1$ always.
1.3. Related Work
Opinion Dynamics
[…]
As in these works, we treat the FJ model as a basis. However, our work is significantly different as it studies a framework where any subset $S\subseteq[n]$ of strategic agents can deviate from their truthful intrinsic opinions, as opposed to studying the evolution of the expressed opinions and their PSNE in the FJ model. In our model, each strategic agent $i\in S$ can only choose a single entry $\bm{s}_{i}^{\prime}$ of the overall deviation $\bm{s}^{\prime}$, but pays a cost based on the resulting equilibrium ($\bm{z}^{\prime}=B\bm{s}^{\prime}$), which depends on the choices of other members of $S$.
Disagreement and Polarization in Social Networks
Motivated by real-world manipulation of social networks in, e.g., the 2016 US election, a recent line of work studies polarization and strategic behavior in opinion dynamics (Gaitonde et al., 2020, 2021; Chen and Rácz, 2021b; Wang and Kleinberg, 2024; Ristache et al., 2024, 2025). Chen and Rácz (2021b) consider a model in which an adversary can control $k\leq n$ nodes’ internal opinions and seeks to maximize polarization at equilibrium. Similarly, Gaitonde et al. (2020) considers a single adversary who can modify intrinsic opinions $\bm{s}$ belonging to an $\ell_{2}$-ball. More recent work also studies modification of agents’ susceptibility parameters $\alpha_{i}$ to alter the median opinion at equilibrium (Ristache et al., 2025). By contrast, we study a setting in which any subset $S\subseteq[n]$ can be strategic. Unlike previous works, these “adversaries” can have conflicting goals in our model.
Manipulation of Dynamic Games.
Opinion dynamics are a widely studied instance of a network game, which is a game played by nodes in a network with payoffs depending on the actions of their neighbors (Kearns et al., 2001; Tardos, 2004). In addition to the manipulation of opinion dynamics, researchers have studied strategic manipulation of financial network formation (Jalan and Chakrabarti, 2024). In the non-network setting, researchers have studied the manipulation of recommendation systmes from a game-theoretic perspective (Ben-Porat and Tennenholtz, 2018), as well as security games (Nguyen et al., 2019), repeated auctions (Kolumbus and Nisan, 2022) and Fisher markers with linear utilities (Kolumbus et al., 2023).
Learning from Strategic Data.
We develop learning algorithms which observe the (possibly manipulated) equilibrium $\bm{z}^{\prime}$ to detect if manipulation occurred, and if so who was responsible. The former problem relates to anomaly detection in networks. Chen and Tsourakakis (2022) develop a hypothesis test to detect such fraud in financial transaction neteworks, by testing if certain subgraphs deviate from Benford’s Law. Similarly, Agarwal et al. (2020) propose a framework based on a $\chi^{2}$-statistic to perform graph similarity search.
The problem of recovering the set of deviators relates to the broader literature of learning from observations of network games. Most works give learning algorithms for games without manipulation (Irfan and Ortiz, 2014; Garg and Jaakkola, 2016; De et al., 2016; Leng et al., 2020; Rossi et al., 2022; Jalan et al., 2024). But our data $\bm{z}^{\prime}$ can be a manipulated equilibrium, which is a strategic sources of data (Zampetakis, 2020). Learning algorithms for strategic sources are known for certain settings such as linear classifiers with small-deviation assumptions (Chen et al., 2020), or binary classifiers in a linear reward model (Harris et al., 2023). When agents can modify their features to fool a known algorithm, even strategy-robust classifiers such as (Hardt et al., 2016) can be inaccurate (Ghalme et al., 2021). Since agents can deviate arbitrarily in our model, we use a robust regression method with guarantees against adversarial corruptions (Bhatia et al., 2015), similar to the learning algorithms in (Kapoor et al., 2019; Russo, 2023). The work of Jalan and Chakrabarti (2024) studies learning from financial networks with strategic manipulations, which is in a similar spirit to our work but differs significantly in the application domain and context.
1.4. Real-world Datasets
To support our results, we use data grounded in practice, which have also been used in previous studies to study polarization and disagreement (cf. Musco et al. (2018b); Chitra and Musco (2020); Wang and Kleinberg (2024); Adamic and Glance (2005)). Specifically, we use Twitter, Reddit, and Political blog networks, summarized in Table 1 summarizes these. Both the Twitter and Reddit datasets are due to Chitra and Musco (2020). The vectors $\bm{s}$ of initial opinions for both are obtained via sentiment analysis and also follow the post-processing of Wang and Kleinberg (2024).
(1) Twitter dataset.
These data correspond to debate over the Delhi legislative assembly elections of 2013. Nodes are Twitter users, and edges refer to user interactions.
(2) Reddit dataset.
These data correspond to political discussion on the r/politics subreddit. Nodes are who posted in the r/politics subreddit, and there is an edge between two users $i,j$ if two subreddits (other than r/politics) exist that both $i,j$ posted on during the given time period.
(3) Political Blogs (Polblogs) dataset.
These data, due to Adamic and Glance (2005), contain opinions from political blogs (liberal and conservative). Edges between blogs were automatically extracted from a crawl of the front page of the blog. Each blog is either liberal – where we assign a value $\bm{s}_{i}=-1$ – or conservative – where we assign $\bm{s}_{i}=+1$.
Network Nodes ($n$) Edges ($m$) Description
Twitter 548 3,638 User interactions during 2013 Delhi elections.
Reddit 556 8,969 User interactions in r/politics subreddit
Polblogs 1,490 16,178 Liberal and conservative blog network
Table 1. Summary of the social network datasets we use.
2. Strategic Opinion Formation
The opinion formation game has two phases. First, strategic agents privately choose a strategic intrinsic opinion according to Equation 3. Second, agents exchange opinions and reach consensus as if they were in the Friedkin-Johnson dynamics, except the strategic opinions are used in place of the true intrinsic opinions.
  1. (1)
    Strategy Phase. Each strategic agent $i\in S$ independently and privately chooses a fictitious strategic opinion $\bm{s}_{i}^{\prime}\in\mathbb{R}$. For honest agents ($i\notin S$) we have $\bm{s}_{i}^{\prime}=\bm{s}_{i}$.
  2. (2)
    Opinion Formation Phase. Reach equilibrium $\bm{z}^{\prime}=B\bm{s}^{\prime}$ as if $\bm{s}^{\prime}$ were the true intrinsic opinions $\bm{s}$.
The network $G$ and the true beliefs $\bm{s}$ determine each agent’s utility. We pose the following problem:
Definition 2.1 (Instrinsic belief lying problem.).
Let $S\subseteq[n]$ be a set of strategic agents. If agent $i\in S$ wants network members to express opinions close to $\bm{s}_{i}$, what choice of $\bm{s}_{i}^{\prime}$ is optimal and minimizes the cost function of Equation 3?
The following theorem characterizes the Nash Equilibria of the Intrinsic Belief Lying Problem.
Theorem 2.2 (Nash Equilibrium).
Let $\mathbb{T}_{i}=(1-\alpha_{i})(B^{T}L_{i}B)+\alpha_{i}(B^{T}\bm{e}_{i}\bm{e}_{i}^{T}B)\in\mathbb{R}^{n\times n}$ and $\bm{y}_{i}=\alpha_{i}B_{ii}\bm{s}_{i}$. The Nash equilibria, if any exist, are given by solutions $s^{\prime}\in\mathbb{R}^{n}$ to the following constrained linear system:
$\forall i\in S:\bm{e}_{i}^{T}\mathbb{T}_{i}\bm{s}^{\prime}$ $=\bm{y}_{i},$
$\forall j\not\in S:\bm{s}_{j}^{\prime}$ $=\bm{s}_{j}.$
Proof.
Consider agent $i\in S$. To calculate the best-response $\bm{s}_{i}^{\prime}$ of $i$ in response to $\bm{s}_{-i}^{\prime}$, we analyze derivatives of its cost function with respect to $\bm{s}^{\prime}$. Since the equilibrium $\bm{z}^{\prime}$ is $\bm{z}^{\prime}=B\bm{s}^{\prime}$, we have:
$c_{i}(\bm{z}^{\prime})$ $=(1-\alpha_{i})\sum_{j\sim i}w_{ij}(\bm{z}_{i}^{\prime}-\bm{z}_{j}^{\prime})^{2}+\alpha_{i}(\bm{z}_{i}^{\prime}-\bm{s}_{i})^{2}$
$c_{i}(\bm{s}^{\prime})$ $=(1-\alpha_{i})\sum_{j\sim i}w_{ij}((\bm{e}_{i}-\bm{e}_{j})^{T}B\bm{s}^{\prime})^{2}+\alpha_{i}(\bm{e}_{i}^{T}(B\bm{s}^{\prime}-\bm{s}))^{2}$
$=(1-\alpha_{i})\sum_{j\sim i}w_{ij}(\bm{s}^{\prime})^{T}(B^{T}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}B)(\bm{s}^{\prime})$
$+\alpha_{i}((\bm{s}^{\prime})^{T}B^{T}e_{i}e_{i}^{T}B\bm{s}^{\prime}-2(\bm{s}^{\prime})^{T}B^{T}e_{i}e_{i}^{T}\bm{s}+\bm{s}^{T}e_{i}e_{i}^{T}\bm{s})$
$\nabla_{\bm{s}^{\prime}}c_{i}(\bm{s}^{\prime})$ $=(1-\alpha_{i})\sum_{j\sim i}w_{ij}2(B^{T}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}B)(\bm{s}^{\prime})+\alpha_{i}(2B^{T}\bm{e}_{i}\bm{e}_{i}^{T}B\bm{s}^{\prime}-2B^{T}\bm{e}_{i}\bm{e}_{i}^{T}\bm{s}),$
$\nabla_{\bm{s}^{\prime}}^{2}c_{i}(\bm{s}^{\prime})$ $=2(1-\alpha_{i})B^{T}\bigg{[}\sum_{j\sim i}w_{ij}2(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}\bigg{]}B+2\alpha_{i}B^{T}\bm{e}_{i}\bm{e}_{i}^{T}B.$
Let $L_{i}\in\mathbb{R}^{n\times n}$ be:
$L_{i}:=\sum_{j\sim i}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}.$
Notice that $L_{i}$ is precisely the Laplacian of the graph when all edges not incident to $i$ are equal to zero. Therefore $L_{i}\succeq 0$. Since $\bm{e}_{i}\bm{e}_{i}^{T}\succeq 0$, the Hessian of $c_{i}$ with respect to $\bm{s}^{\prime}$ is PSD. In particular, its $(i,i)$ entry is non-negative, so $\frac{\partial^{2}c_{i}(\bm{s}^{\prime})}{\partial(\bm{s}_{i}^{\prime})^{2}}\geq 0$, and hence the optimal $\bm{s}_{i}^{\prime}$ is at the critical point. This is given as:
00 0 $=\frac{1}{2}\frac{\partial}{\partial\bm{s}_{i}^{\prime}}c_{i}(\bm{s}^{\prime})$
$=\bm{e}_{i}^{T}(1-\alpha_{i})B^{T}\bigg{[}\sum_{j\sim i}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}\bigg{]}B\bm{s}^{\prime}+\bm{e}_{i}^{T}\alpha_{i}(B^{T}\bm{e}_{i}\bm{e}_{i}^{T}Bs^{\prime}-B^{T}\bm{e}_{i}\bm{e}_{i}^{T}\bm{s})$
$=(1-\alpha_{i})\bm{e}_{i}^{T}B^{T}L_{i}B\bm{s}^{\prime}+\bm{e}_{i}^{T}\alpha_{i}(B^{T}\bm{e}_{i}\bm{e}_{i}^{T}B\bm{s}^{\prime}-B^{T}\bm{e}_{i}\bm{e}_{i}^{T}\bm{s}).$
The above display gives the solution for $\bm{s}_{i}^{\prime}$ in terms of all entries of $s^{\prime}$. Assembling the critical points into a linear system, we obtain precisely that for all $i\in S$, $\bm{e}_{i}^{T}\mathbb{T}_{i}\bm{s}^{\prime}=\bm{y}_{i}$. Since $\bm{s}_{j}^{\prime}=\bm{s}_{j}$ for $j\not\in S$, the overall linear system describes the Nash equilibria. ∎
To illustrate the Theorem, we consider a toy example.
Example 2.3 (Two-Node Graph).
Consider a graph with 2 nodes and one edge with weight $w>0$. We set $\alpha_{1}=\alpha_{2}=0.5$ for simplicity. Suppose that both agents deviate, i.e., $S=[2]$. Then, we can calculate $B$ to be
$B=\frac{1}{2w+1}\begin{pmatrix}w+1&w\\ w&w+1\end{pmatrix}$
and
(9) $\bm{z}_{0}^{\prime}=\frac{(w+1)\bm{s}_{0}^{\prime}+w\bm{s}_{1}^{\prime}}{2w+1},\quad\bm{z}_{1}^{\prime}=\frac{w\bm{s}_{0}^{\prime}+(w+1)\bm{s}_{1}^{\prime}}{2w+1},$
yielding the two cost functions
$c_{0}(\bm{s}^{\prime})$ $=\frac{1}{2}w\left(\frac{\bm{s}_{0}^{\prime}-\bm{s}_{1}^{\prime}}{2w+1}\right)^{2}+\frac{1}{2}\left(\frac{(w+1)\bm{s}_{0}^{\prime}+w\bm{s}_{1}^{\prime}}{2w+1}-\bm{s}_{0}\right)^{2}$
$c_{1}(\bm{s}^{\prime})$ $=\frac{1}{2}w\left(\frac{\bm{s}_{0}^{\prime}-\bm{s}_{1}^{\prime}}{2w+1}\right)^{2}+\frac{1}{2}\left(\frac{(w+1)\bm{s}_{1}^{\prime}+w\bm{s}_{0}^{\prime}}{2w+1}-\bm{s}_{1}\right)^{2}.$
Taking the first order conditions $\frac{\partial c_{0}}{\partial\bm{s}_{0}^{\prime}}=0$ and $\frac{\partial c_{1}}{\partial\bm{s}_{1}^{\prime}}=0$ we get a linear system whose solutions are:
$\bm{s}_{0}^{\prime}$ $=\frac{w^{2}(s_{0}-s_{1})+(3w+1)\bm{s}_{0}}{3w+1},\quad\bm{s}_{1}^{\prime}=\frac{w^{2}(\bm{s}_{0}-\bm{s}_{1})+(3w+1)s_{1}}{3w+1}.$
Replacing these values back to the costs we get that
$\forall i:c_{i}(s_{0}^{\prime},s_{1}^{\prime})=\frac{1}{2}\frac{w(w^{2}+3w-1)(\bm{s}_{0}-\bm{s}_{1})^{2}}{9w^{2}+6w+1},$
On the other hand, if all agents are honest, then the cost for each is:
$\forall i:c_{i}(\bm{s}_{0},\bm{s}_{1})=\frac{1}{2}\frac{w(w+1)(\bm{s}_{0}-\bm{s}_{1})^{2}}{(2w+1)^{2}}.$
and the ratio of the two costs is at least $\max\{1,w/3\}$.
Next, we discuss some consequences of Theorem 2.2. First, we characterize $\bm{s}^{\prime}$ as the solution to a linear system.
Corollary 2.4.
Let $T\in\mathbb{R}^{|S|\times n}$ have rows $\{\bm{e}_{i}^{T}\mathbb{T}_{i}\}_{i\in S}$ given by Theorem 2.2. Let $\tilde{T}\in\mathbb{R}^{|S|\times|S|}$ be the submatrix of $T$ selecting columns belonging to $S$. Let $\bm{y}\in\mathbb{R}^{|S|}$ have entries $\bm{y}_{i}=\alpha_{i}B_{ii}\bm{s}_{i}$ as above. Let $\tilde{\bm{y}}=\bm{y}-\sum_{j\not\in S}\bm{s}_{j}T\bm{e}_{j}$. Then the set of Nash equilibria, if any exist, are given by the solutions to the unconstrained linear system
(10) $\tilde{T}\bm{x}=\tilde{\bm{y}}.$
The resulting opinions vector $\bm{s}^{\prime}$ is given by $\bm{s}_{i}^{\prime}=\bm{x}_{i}$ if $i\in S$ and $\bm{s}_{i}^{\prime}=\bm{s}_{i}$ otherwise.
Thus, in a Nash equilibrium, every strategic agent solves their corresponding equation given by Equation 10. The explicit characterization of equilibria also implies that Nash equilibria cannot be mixed.
Corollary 2.5 (Pure Strategy Nash Equilibria).
The Nash equilibrium corresponds to solving the system of $|S|$ linear equations in the scalars $\{\bm{s}_{i}^{\prime}|i\in S\}$ given by Equation 10. Also, all Nash equilibria are pure-strategy Nash equilibria.
Figure 2. Plot of truthful intrinsic opinions ($s$) and strategic opinions ($s^{\prime}$), and truthful public opinions ($z$) compared to the strategic public opinions ($z^{\prime}$) for the nodes belonging to $S$. $S$ is taken to be the top-50% in terms of their eigenvector centrality. In both cases we have taken $\alpha_{i}\in\{0.25,0.5\}$ for all nodes. We fit a linear regression between $s^{\prime}$ and $s$ (resp. between $z$ and $z^{\prime}$). We report the effect size $\theta$ which corresponds to the slope of the linear regression and the $P$-value with respect to the null hypothesis ($\theta=0$). ∗∗∗ stands for $P<0.001$, ∗∗ stands for $P<0.01$ and stands for $P<0.05$.
Figure 3. Strategic misreports for the Polblogs dataset where $S$ is taken to be the top-50% of the agents in terms of their eigenvector centralities. The nodes are labeled either as liberal ($\bm{s}_{i}=-1$) or conservative ($\bm{s}_{i}=+1$), and we consider the nodes that change their beliefs as the nodes for which $\bm{z}_{i}^{\prime}$ and $\bm{z}_{i}$ do not have the same sign. In the scatterplots (a), (c), (d), (e), the shape of each point indicates whether that user changed belief or not, and the color indicates their true (intrinsic) opinion. Overall, we discover a higher amount of change among liberal blogs compared to conservative ones (panel (b)). Additionally, we report the truthful/strategic public opinion as a function of the logarithm of the eigenvector centrality $\bm{\pi}_{i}$ (cf. panels (c, d)) for each node, as well as the absolute change $|\bm{z}_{i}^{\prime}-\bm{z}_{i}|$ (cf. panel (e)). We fit a regression model, and we detect significant effects (${}^{***}:P<0.001,\;^{**}:P<0.01,\;^{*}:P<0.05$; effects denoted by $\theta$) of the logarithm of the centrality to the truthful equilibrium $\bm{z}$, the strategic equilibrium $\bm{z}^{\prime}$, and the change $|\bm{z}^{\prime}-\bm{z}|$, revealing the structure of a power law. Finally, we observe that relative changes are more dispersed along liberal sources compared to conservative sources (cf. panel (f)).
Optimal Deviation for One Agent and All Agents
Assuming that we have one strategic agent, what is the change in their opinion? We can show that the new opinion is a scalar multiple of the initial opinion plus a bias term, where neither the scalar multiple nor the bias term can be zero.
Corollary 2.6 (Deviation for One Agent).
Let $S=\{i\}$. Then, $\bm{s}_{i}^{\prime}=\theta_{i}\bm{s}_{i}+\beta_{i}$ where
$\theta_{i}$ $=\frac{\alpha_{i}B_{ii}}{(1-\alpha_{i})\sum_{i\sim j}w_{ij}(B_{ii}-B_{ij})^{2}+\alpha_{i}B_{ii}^{2}}>0,$
$\beta_{i}$ $=-\frac{\alpha_{i}\sum_{j\neq i}B_{ij}s_{j}}{(1-\alpha_{i})\sum_{i\sim j}w_{ij}(B_{ii}-B_{ij})^{2}+\alpha_{i}B_{ii}^{2}}.$
Similarly, we can relate the maximum deviation of $\bm{s}^{\prime}$ from $\bm{s}$ in the other extreme case, i.e., when all agents are deviating ($S=[n]$).
Corollary 2.7.
When all agents are deviating ($S=[n]$), and $\alpha_{i}=\alpha$, then $\bm{s}^{\prime}$ satisfies:
$\frac{\|\bm{s}^{\prime}\|_{2}}{\|\bm{s}\|_{2}}\leq\frac{\lambda_{n}+\tilde{\alpha}}{\tilde{\alpha}}.$
Proof of Corollary 2.7.
When all agents are deviating, it is straightforward to show that $\tilde{T}=\tilde{\alpha}B$ with minimum eigenvalue $\tilde{\alpha}^{2}/(\lambda_{n}+\tilde{\alpha})>0$. Thus $\tilde{T}$ is invertible, and therefore $\bm{s}^{\prime}=\frac{1}{\tilde{\alpha}}B^{-1}\widetilde{\textrm{diag}(B)}\bm{s}$, where $\widetilde{\textrm{diag}(B)}$ is a diagonal matrix with entries $B_{ii}$. Then
$\|\bm{s}^{\prime}\|_{2}$ $\leq\frac{1}{\tilde{\alpha}}\|B^{-1}\|_{2}\|\widetilde{\textrm{diag}(B)}\|_{2}\|\bm{s}\|_{2}$
$=\left(\max_{i}B_{ii}\right)\left(\max_{i}\frac{\lambda_{i}+\tilde{\alpha}}{\tilde{\alpha}}\right)\|\bm{s}\|_{2}$
$\leq\frac{\tilde{\alpha}}{\lambda_{1}+\tilde{\alpha}}\frac{\lambda_{n}+\tilde{\alpha}}{\alpha}\|\bm{s}\|_{2}$
$=\frac{\lambda_{n}+\tilde{\alpha}}{\tilde{\alpha}}\|\bm{s}\|_{2}.$
The proof of Corollary 2.7 shows that the adjusted susceptibility ($\tilde{\alpha}$) and the maximum eigenvalue of the Laplacian ($\lambda_{n}$) are responsible for changes in the norm of $\bm{s}^{\prime}$. From classic spectral graph theory, we know that $\lambda_{n}=\Theta(d_{\textup{max}})$ where $d_{\textup{max}}$ is the maximum degree of the graph; therefore, graphs with a lower maximum degree experience smaller distortions. Also, regarding the susceptibility to persuasion, the distortion becomes $1+o(1)$ as long as $\tilde{\alpha}=\omega(d_{\textup{max}})$.
Equilibria for real-world datasets. Next, we discuss the results of experiments simulating the strategically manipulated equilibria for our real-world datasets.
Figure 4. Polarization ratio ($\mathcal{P}(z^{\prime})/\mathcal{P}(z)$), disagreement ratio ($\mathcal{D}(z^{\prime})/\mathcal{D}(z)$), and price of misreporting ($C(z^{\prime})/C(z)$) for the three datasets for varying susceptibility to persuasion values. We have set all susceptibilities $\alpha_{i}$ to the same value $\alpha$. The Twitter dataset has the largest variation in all three ratios compared to the others. $S$ is taken to be the top-50% nodes in terms of their eigenvector centrality.
Effect of Susceptibility to Persuasion in Real-world Data
Regarding real-world data, Figure 2 shows the relationship between the truthful opinions ($\bm{s}$ and $\bm{z}$) and the strategic ones ($\bm{s}^{\prime}$ and $\bm{z}^{\prime}$) for the datasets, along with the corresponding correlation coefficient $R^{2}$, assuming that $S$ consists of the top-50% nodes in terms of their eigenvector centrality, for susceptibility parameters set to $\alpha_{i}=0.5$ (equal self-persuasion and persuasion due to others) and $\alpha_{i}=0.25$ (higher persuasion due to others).
Regarding the public opinions, even though in the Reddit dataset, the strategic opinions seem to be correlated with the truthful ones ($R^{2}=0.78$ for $\alpha_{i}=0.25$ and $R^{2}=0.94$ for $\alpha_{i}=0.5$ respectively), in the Twitter dataset, we do not get the same result (i.e., $R^{2}<0.25$). Finally, in the Polblogs dataset, the situation is somewhere in the middle; when $\alpha_{i}=0.25$ we get a low $R^{2}$ ($R^{2}=0.18$) where for $\alpha_{i}=0.5$ we get a high $R^{2}$ ($R^{2}=0.74$). Additionally, in all cases except Twitter, we get that the effect is significant ($P<0.01$).
Regarding the relationship between the intrinsic opinions, we do not detect any significant effect in most cases except Reddit with $\alpha_{i}=0.5$ ($P<0.01$) and Twitter with $\alpha_{i}=0.5$ ($P<0.05$).
Asymmetric Effects of Strategic Behavior on Liberals and Conservatives.
Figure 3 analyzes the opinions of the strategic set $S$ on the Polblogs dataset. Specifically, we find that larger changes in sentiment happen across liberal outlets compared to conservative ones. Additionally, the changes in the truthful/strategic opinions are related to the eigenvector centrality $\bm{\pi}_{i}$ as a power law, i.e., $\bm{z}_{i}^{\prime}\propto\bm{\pi}_{i}^{\theta}$ ($P<0.001$; linear regression between the log centralities $\log\bm{\pi}_{i}$ and $\bm{z}_{i}^{\prime}$). The same finding holds for $|\bm{z}_{i}^{\prime}-\bm{z}_{i}|$ and $\bm{z}_{i}$.
At this point, one may wonder whether the eigenvector centrality really influences the strategic opinions $\bm{z}_{i}^{\prime}$ for $i\in S$. Our answer is negative. We repeat the same experiment with the Twitter and Reddit datasets, where we find no effects ($P>0.1$; linear regression between the log centralities $\log\bm{\pi}_{i}$ and $\bm{z}_{i}^{\prime}$). Due to space limitations, the corresponding figures are deferred to Appendix C.
Polarization and Disagreement.
Figure 4 shows how the polarization, disagreement, and cost change as a function of the susceptibility parameter $\alpha_{i}$. Except for $\alpha_{i}\approx 0.3$, the polarization ratio, disagreement ratio, and the price of misreporting experience a downward trend as $\alpha_{i}$ increases. This indicates that as as users prioritize their own opinions more than their neighbors, they are less susceptible to strategic manipulation.
Effect of the number of deviators ($|S|$)
Next, we study the effect of the number of deviators, which corresponds to $|S|$, on the changes in polarization, disagreement, and the total cost (through the price of misreporting). Figure 5 shows how the polarization and disagreement when $S$ consists of the top-1-10% most central agents with respect to eigenvector centrality. We show that even if only 1% of agents are strategic, this can impact consensus by several orders of magnitude.
Figure 5. Polarization ratio ($\mathcal{P}(\bm{z}^{\prime})/\mathcal{P}(\bm{z})$), disagreement ratio ($\mathcal{D}(\bm{z}^{\prime})/\mathcal{D}(\bm{z})$), and price of misreporting ($C(\bm{z}^{\prime})/C(\bm{z})$) for the three datasets for varying the size of $|S|$. The size of $|S|$ corresponds to the top $p$ percent of the actors ($|S|=\lceil pn\rceil$) based on their eigenvector centrality (in decreasing order), for $p\in[0.01,0.1]$. The susceptibility parameter is set to $\alpha_{i}=0.5$.
3. Price of Misreporting
In Section 2, we saw that strategic manipulation can substantially affect network outcomes via the Polarization Ratio and Disruption Ratio. We now give an upper bound for the Price of Misreporting (Eq. (8)), which is the analogue of the Price of Anarchy in our setting. The $\mathsf{PoM}$ measures the total cost paid by agents under the corrupted equilibrium $\bm{z}^{\prime}$, versus the total cost under the non-corrupted $\bm{z}$. Since the cost captures an agent’s deviation from her truthful intrinsic opinion as well as her deviation from the expressed opinions of her neighbors, it is a natural measure of the network’s discord at equilibrium.
Theorem 3.1 shows that the PoM is small when the spectral radius of the Laplacian is small, and when agents are somewhat susceptible to their neigbhors ($\alpha\not\to 0$). Note that the spectral radius can be replaced by a degree bound: if $d_{\textup{max}}$ is the maximum degree of the graph, then $\lambda_{n}\leq 2d_{\textup{max}}$. So the PoM is small if the maximum degree is small.
Theorem 3.1.
Suppose all agents deviate ($S=[n]$) and there exists $\alpha$ such that $\alpha_{i}=\alpha$ for all $i$. Let $\tilde{\alpha}=\alpha/(1-\alpha)$, and $\lambda_{n}$ be the spectral radius of the Laplacian. Then the price of misreporting is bounded as:
$\mathsf{PoM}\leq\frac{(\lambda_{n}+4\tilde{\alpha})(\lambda_{n}+\tilde{\alpha})^{2}}{\tilde{\alpha}^{5}}=O\left(\max\left\{\frac{\lambda_{n}}{\tilde{\alpha}^{5}},\frac{1}{\tilde{\alpha}^{2}}\right\}\right).$
Proof of Theorem 3.1..
First, we set $\tilde{\alpha}=\alpha/(1-\alpha)$. By substituting $\bm{z}=B\bm{s}$ we can show by straightforward algebra that $C(\bm{z})/(1-\alpha)=\bm{s}^{T}Q\bm{s}$ where $Q\succeq 0$ with
(11) $Q=BLB+\tilde{\alpha}(I-2B+B^{2})$
Since $Q\succeq 0$, it has eigendecomposition $Q=U\Lambda_{Q}U^{T}$. Moreover, $U$ is precisely the matrix of eigenvectors for the Laplacian. The eigenvalues of $Q$ can be shown to be $\tilde{\alpha}^{2}/(\lambda_{i}+\tilde{\alpha})$. Therefore, $C(z)=\bm{s}^{T}Q\bm{s}\geq\frac{\tilde{\alpha}^{2}}{\lambda_{n}+\tilde{\alpha}}\|\bm{s}\|_{2}^{2}$.
Next, let $\textrm{diag}(B)$ be the diagonal matrix with entries $B_{ii}$ and $\widetilde{\textrm{diag}(B)}$ be as in Corollary 2.7. In the proof of Corollary 2.7, we show that $\bm{s}^{\prime}=(1/\tilde{\alpha})B^{-1}\widetilde{\textrm{diag}(B)}\bm{s}$ and $\bm{z}^{\prime}=(1/\tilde{\alpha})\widetilde{\textrm{diag}(B)}\bm{s}$, which similarly implies (after algebraic operations) that $C(\bm{z}^{\prime})/(1-\alpha)=\bm{s}^{T}Q^{\prime}\bm{s}$ where:
$Q^{\prime}:=\frac{1}{\tilde{\alpha}^{2}}\widetilde{\textrm{diag}(B)}L\widetilde{\textrm{diag}(B)}+\frac{1}{\tilde{\alpha}}B^{-1}\left(\widetilde{\textrm{diag}(B)}\right)^{2}B^{-1}-2\frac{1}{\tilde{\alpha}}B^{-1}\left(\widetilde{\textrm{diag}(B)}\right)^{2}+\frac{1}{\tilde{\alpha}}\left(\widetilde{\textrm{diag}(B)}\right)^{2}$
Note that $Q^{\prime}$ cannot be diagonalized since, in general, $\widetilde{\textrm{diag}(B)}$ has a different eigenbasis than $L$. However, we note that:
(12) $\|\widetilde{\textrm{diag}(B)}\|_{2}$ $=\max_{i}B_{ii}\leq\|B\|_{2}=1,$
(13) $\|B^{-1}\|_{2}$ $=\max_{i}\frac{\lambda_{i}+\tilde{\alpha}}{\tilde{\alpha}}=\frac{\lambda_{n}+\tilde{\alpha}}{\tilde{\alpha}}.$
By the triangle inequality, the Cauchy-Schwarz inequality, and Equations 12 and 13, we have that:
$\|Q^{\prime}\|_{2}$ $\leq\frac{1}{\tilde{\alpha}^{2}}\|L\|_{2}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}+\frac{1}{\tilde{\alpha}}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}\|B^{-1}\|_{2}^{2}+\frac{2}{\tilde{\alpha}}\|B^{-1}\|_{2}+\frac{1}{\tilde{\alpha}}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}$
$\leq\frac{(\lambda_{n}+4\tilde{\alpha})(\lambda_{n}+\tilde{\alpha})}{\tilde{\alpha}^{3}}.$
Therefore $C(\bm{z}^{\prime})/(1-\alpha)\leq\frac{(\lambda_{n}+4\tilde{\alpha})(\lambda_{n}+\tilde{\alpha})}{\tilde{\alpha}^{3}}\|\bm{s}\|_{2}^{2}$. Hence,
(14) $\frac{C(\bm{z}^{\prime})}{C(\bm{z})}\leq\frac{(\lambda_{n}+4\tilde{\alpha})(\lambda_{n}+\tilde{\alpha})^{2}}{\tilde{\alpha}^{5}}.$
Finally, we can simplify:
(15) $\frac{(\lambda_{n}+4\tilde{\alpha})(\lambda_{n}+\tilde{\alpha})^{2}}{\tilde{\alpha}^{5}}\leq\frac{64(\lambda_{n}+\tilde{\alpha})^{3}}{\tilde{\alpha}^{5}}\leq\frac{128(\max\{\lambda_{n},\tilde{\alpha}\})^{3}}{\tilde{\alpha}^{5}}.$
From Theorem 3.1, we can show that the upper bound is minimized when $\lambda_{n}=\Theta(\tilde{\alpha}^{3})$ and has a value of $O(1/\tilde{\alpha}^{2})$. As we noted, Theorem 3.1 can be written with $d_{\textup{max}}$ in the place of $\lambda_{n}$ as well.
Next, we give an easy generalization to the case of differing susceptibility.
Corollary 3.2 (Price of Misreporting for Heterogeneous Susceptibility).
If the $\alpha_{i}$ are differing, let $\alpha_{\min}=\min_{i}\alpha_{i}$ and $\alpha_{\max}=\max_{j}\alpha_{j}$. Define $\tilde{\alpha}_{\min}=\frac{\alpha_{\min}}{1-\alpha_{\max}},\tilde{\alpha}_{\max}=\frac{\alpha_{\max}}{1-\alpha_{\min}}$. The Price of Misreporting is bounded as:
$\mathsf{PoM}\leq\frac{1-\alpha_{\min}}{1-\alpha_{\max}}\frac{(\lambda_{n}+4\tilde{\alpha}_{\max})(\lambda_{n}+\tilde{\alpha}_{\max})^{2}}{\tilde{\alpha}_{\min}}.$
Proof.
Note that $C(\bm{z}^{\prime})\leq(1-\alpha_{\min})(\bm{z}^{\prime})^{T}L\bm{z}^{\prime}+\alpha_{\max}\|\bm{z}^{\prime}-\bm{s}\|_{2}^{2}=\overline{C}(\bm{z}^{\prime})$, and $C(\bm{z})\geq(1-\alpha_{\max})\bm{z}^{T}L\bm{z}+\alpha_{\min}\|\bm{z}-\bm{s}\|_{2}^{2}=\underline{C}(\bm{z})$ where $\alpha_{\min}=\min_{i\in[n]}\alpha_{i}$, and $\alpha_{\max}=\max_{i\in[n]}\alpha_{i}$. Then, the same analysis of Theorem 3.1 can be applied, since $\overline{C}(\bm{z}^{\prime})/\underline{C}(\bm{z})\geq C(\bm{z}^{\prime})/C(\bm{z})$. ∎
Finally, we discuss how one may generalize Theorem 3.1 to the case where some agents are honest.
Towards fine-grained $\mathsf{PoM}$ guarantees.
Figure 4 shows that the PoM is not monotonic in $\lvert S\rvert$. As the number of strategic agents grows, the PoM can fall or grow, depending on the choice of $S$, network parameters, and so on. Therefore, we would like to give a version of Theorem 3.1 for any set of strategic agents $S\subset[n]$, not just the case of $S=[n]$. However, proving such a bound would require analyzing $S\times S$ principal submatrices of $B,L$ to obtain characterizations of the cost at the corrupted equilibrium $\bm{z}^{\prime}$. In particular, we would require a restricted invertibility estimate to prove the analogue of Eq. (13). To our knowledge, the best such estimates (Marcus et al., 2022) are too lossy when $n-\lvert S\rvert$ is large. We leave this question to future work.
4. Learning from Network Outcomes
To mitigate the effects of strategic behavior, a platform must understand whether manipulation has occurred, and who the strategic actors are. In this section, we give computationally efficient methods to do so based on knowledge of the network edges and observing the corrupted equilibrium $\bm{z}^{\prime}$. The latter can be found, for example, by performing sentiment analysis on the users’ posts.
4.1. Detecting Manipulation with a Hypothesis Test
In many real-world networks, the distribution of truthful opinions follows a Gaussian distribution (Figure 6). Given estimates $(\widehat{L},\widehat{A})$ for the graph Laplacian and susceptibility matrix, the platform can observe the corrupted equilibrium $\bm{z}^{\prime}$ and solve for the strategic opinions $\bm{s}^{\prime}$ via:
(16) $\widehat{\bm{s}^{\prime}}:=\widehat{A}^{-1}((I-\widehat{A})\widehat{L}+\widehat{A})\bm{z}^{\prime}.$
We propose that the platform perform a one-sample $t$-test with the entries $\bm{s}^{\prime}$, with a population mean $\mu_{0}\in\mathbb{R}$ based on e.g. historical data. Under the null hypothesis in which no manipulation has occurred, $\bm{s}^{\prime}=\bm{s}$, so the test should fail to reject the null hypothesis. However, when agents $S\subset[n]$ deviate, then $\bm{s}_{i}^{\prime}\neq\bm{s}_{i}$ for $i\in S$, so the test should reject the null for large enough deviations. The test is simple, and described in Algorithm 1. Figure 6 shows the results of the test for varying choices of $S$. We see that at significance level $0.05$, the test has low Type I error, as it will return “No Manipulation” when $S=\emptyset$, and low Type II error as it will return “Manipulation” when $S\neq\emptyset$.
Input: Estimated graph information $\widehat{L}$ and $\widehat{A}$, observed equilibrium $\bm{z}^{\prime}$
Output: “Manipulation” or “No Manipulation”
Observe corrupted equilibrium $\bm{z}^{\prime}$.
Solve for $\widehat{\bm{s}^{\prime}}$ (Eq. (16)).
Perform one-sample $t$ test on the entries of $\widehat{\bm{s}^{\prime}}$ with a population mean $\mu_{0}$ under the null hypothesis.
return If the $t$-test rejects, return “Manipulation.” Otherwise, return “No Manipulation.”
ALGORITHM 1 Learning from Misreporting Equilibrium with Hypothesis Testing
For the Political Blogs dataset, intrinsic opinions belong to {±$\{\pm 1\}$, so the null hypothesis should be a biased Rademacher distribution. In this case, one should use a $\chi^{2}$-test, as in  Agarwal et al. (2020); Chen and Tsourakakis (2022).
Figure 6. The true opinions for Twitter (left) and Reddit (middle) both follow a normal distribution. When simulating strategic manipulation with random choices of $S$ (right), the detection test (Algorithm 1) has no Type I or Type II error at significance level $0.05$. The tester uses $\widehat{L}=L,\widehat{A}=A=\frac{1}{2}I$, and $\mu_{0}$ equal to the mean of the true intrinsic opinions. Shaded regions are $95\%$ confidence intervals for $p$-values of the test across $5$ independent runs.
4.2. Learning the Strategic Actors with Robust Regression
The algorithm described in the previous section (Algorithm 1) can be used to detect whether there exists manipulation in the network. However, the set $S$ is unknown, and therefore the platform cannot target the deviators to perform interventions to mitigate strategic behavior.
It is, therefore, essential for the platform to be able to identify the set of deviators $S$, in case the platform needs to take regulatory actions. While at first, it may seem that finding the set of deviators $S$ is a hard task, it turns out that under mild assumptions on the intrinsic opinion formation process, we can learn the set of deviators $S$ from observing the strategically corrupted equilibrium $\bm{z}^{\prime}$ via Algorithm 2 in polynomial time, described in Algorithm 2. Our algorithm is based on robust regression leveraging the Torrent algorithm developed by Bhatia et al. (2015) and requires access to a node embedding matrix $X\in\mathbb{R}^{n\times d}$, and the size $|S|$ of the set of deviators.
Input: Features $X\in\mathbb{R}^{n\times d}$, graph information $L$ and $A$, observed equilibrium $\bm{z}^{\prime}$, set size $|S|$
Output: Set of strategic agents $\hat{S}$, estimated intrisic beliefs $\hat{\bm{s}}$.
$\widehat{\bm{s}^{\prime}}\leftarrow A^{-1}((I-A)L+A)\bm{z}^{\prime}$;
$\hat{v}\leftarrow$ Robust Regression (Torrent) with design matrix $X$, response vector $\widehat{\bm{s}^{\prime}}$;
$\hat{\bm{s}}\leftarrow X\hat{\bm{v}}$;
$\text{diffs}\leftarrow\lvert\hat{\bm{s}}-\widehat{\bm{s}^{\prime}}\rvert$;
$\hat{S}\leftarrow\text{indices of top }k\text{ largest values in diffs}$;
return $\hat{\bm{s}}\in\mathbb{R}^{n},\hat{S}\subseteq[n]$
ALGORITHM 2 Learning from Misreporting Equilibrium
The key idea of Algorithm 2 is that if the size of the strategic set $S$ is sufficiently small, in general, $|S|\leq Cn$ for some small constant $C$, then we can view the misreported intrinsic opinions $\bm{s}^{\prime}$ as a perturbation of the truthful opinion vector $\bm{s}$, and then use a robust regression algorithm to estimate $\bm{s}$. We assume that the embedding matrix $X\in\mathbb{R}^{n\times d}$ determines intrinsic beliefs: for example, demography, geographic location, etc. Node-level features can be learned by a variety of methods, such as spectral embeddings on the graph Laplacian or graph neural networks. Previous works have used the framework of combining a robust estimator with model-specific information to learn from “strategic sources” of data, such as in bandits (Kapoor et al., 2019), controls (Russo, 2023), and network formation games (Jalan and Chakrabarti, 2024).
In the sequel, we give the precise technical condition of the features required for robust regression to work (Bhatia et al., 2015), which is based on conditions on the minimum and maximum eigenvalues of the correlation matrix determined by the features corresponding to agents in $S$. Specifically, for a matrix $X\in\mathbb{R}^{n\times d}$ with $n$ samples in $\mathbb{R}^{d}$ and $S\subset[n]$ let $X_{S}\in\mathbb{R}^{\lvert S\rvert\times d}$ select rows in $S$. Note that $\lambda_{\min}(\cdot),\lambda_{\max}(\cdot)$ are the min/max eigenvalues respectively.
Definition 4.1 (SSC and SSS Conditions).
Let $\gamma\in(0,1)$. The features matrix $X\in\mathbb{R}^{n\times d}$ satisfies the Subset Strong Convexity Property at level $1-\gamma$ and Subset Strong Smoothness Property at level $\gamma$ with constants $\xi_{1-\gamma},\Xi_{\gamma}$ respectively if:
$\xi_{1-\gamma}$ $\leq\min\limits_{S\subset[n]:\lvert S\rvert=(1-\gamma)n}\lambda_{\min}(X_{S}^{T}X_{S}),$
$\Xi_{\gamma}$ $\geq\max\limits_{S\subset[n]:\lvert S\rvert=\gamma n}\lambda_{\max}(X_{S}^{T}X_{S}).$
We give our guarantee for Algorithm 2.
Proposition 4.2.
Let $X$ be as in Algorithm 2, and suppose that $X\bm{v}=\bm{s}$ for some $\bm{v}\in\mathbb{R}^{d}$, and that $X$ satisfies the SSC condition at level $1-\gamma$ with constant $\xi_{1-\gamma}$, and SSS condition at level $\gamma$ with constant $\Xi_{\gamma}$ (Definition 4.1). Then, there exist absolute constants $C,C^{\prime}>0$ such that if $\lvert S\rvert\leq Cn$ and $14\frac{\sqrt{\Xi_{\gamma}}}{\sqrt{\xi_{1-\gamma}}}<$ 1, Algorithm 2 returns $\hat{\bm{s}}$ such that:
$\|\hat{\bm{s}}-\bm{s}\|_{2}\leq\|X\|_{2}n^{-\omega(1)},$
using $T=C^{\prime}(\log n)^{2}$ iterations of Torrent for the Robust Regression step. Moreover, if for all $j\in S$ we have $\lvert\bm{s}_{j}-\bm{s}_{j}^{\prime}\rvert\gg\|X\|_{2}n^{-\omega(1)}$, then $\hat{S}=S$.
For the proof, we first state the technical result of Bhatia et al. (2015) that we require.
Theorem 4.3 (Bhatia et al. (2015)).
Let $X\in\mathbb{R}^{n\times d}$ be a design matrix and $C>0$ an absolute constant. Let $\bm{b}\in\{0,1\}^{n}$ be a corruption vector with $\|\bm{b}\|_{0}\leq\alpha n$, for $\alpha\leq C$. Let $\bm{y}=X\bm{w}^{*}+\bm{r}$ be the observed responses, and $\gamma\geq\alpha$ be the active set threshold given to the Algorithm 2 of Bhatia et al. (2015). Suppose $X$ satisfies the SSC property at level $1-\gamma$ and SSS property at level $\gamma$, with constants $\xi_{1-\gamma}$ and and $\Xi_{\gamma}$ respectively. If the data $(X,\bm{y})$ are such that $\frac{4\sqrt{\Xi_{\gamma}}}{\sqrt{\xi_{1-\gamma}}}<1$, then after $t$ iterations, Algorithm 2 of Bhatia et al. (2015) with active set threshold $\zeta\geq\gamma$ obtains a solution $\bm{w}^{T}\in\mathbb{R}^{d}$ such that, for large enough $n$,
$\|\bm{w}^{T}-\bm{w}^{*}\|_{2}$ $\leq\frac{\|\bm{r}\|_{2}}{\sqrt{n}}\exp(-cT).$
We are ready to prove Proposition 4.2.
Proof of Proposition 4.2.
Let $\widehat{\bm{s}^{\prime}}$ be as in Algorithm 2, and $\bm{y}=\widehat{\bm{s}^{\prime}}$. Notice $\bm{y}=\bm{s}+(\bm{s}^{\prime}-\bm{s})+(\widehat{\bm{s}^{\prime}}-\bm{s}^{\prime})$. Let $\bm{w}:=(\bm{s}^{\prime}-\bm{s})$ be the corruption vector due to strategic negotiations and $\bm{r}=(\widehat{\bm{s}^{\prime}}-\bm{s}^{\prime})$ be the residual vector due to least-squares regression. We claim that $\bm{r}=\bm{0}$, because $A^{-1}$ is full rank and $((I-A)L+A)$ is full rank, so $\widehat{\bm{s}^{\prime}}=A^{-1}((I-A)L+A)\bm{z}^{\prime}=\bm{s}^{\prime}$.
Next, we apply Theorem  4.3. Notice that $\|\bm{w}\|_{0}\leq Cn$ by assumption. Moreover, $X$ satisfies the SSC and SSS coniditons. Therefore, after $T$ iterations, Algorithm 2 obtains $\hat{\bm{s}}$ such that:
$\|\hat{\bm{v}}-\bm{v}\|_{2}\leq\frac{\exp(-cT)}{\sqrt{n}}\|\bm{s}^{\prime}-\bm{s}\|_{2}.$
Therefore, letting $T=C^{\prime}(\log n)^{2}$ for large enough constant $C^{\prime}>0$, we see that $\|\hat{\bm{v}}-\bm{v}\|_{2}\leq O(n^{-\omega(1)})$. Hence $\|\hat{\bm{s}}-\bm{s}\|_{2}=\|X\hat{\bm{v}}-X\bm{v}\|_{2}\leq\|X\|_{2}\cdot n^{-\omega(1)}$. Now, let $\bm{u}=\hat{\bm{s}}-\bm{s}^{\prime}$. If $i\in[n]\setminus S$, then $\lvert\bm{u}_{i}\rvert\leq\|X\|_{2}\cdot n^{-\omega(1)}$. On the other hand for $j\in S$, $\lvert\bm{u}_{j}\rvert\geq\lvert\bm{s}_{j}-\bm{s}_{j}^{\prime}\rvert-\|X\|_{2}n^{-\omega(1)}$. Therefore the top-$|S|$ entries of $\bm{u}$ recover $S$. ∎
It is interesting to investigate what an upper bound on the size of $S$ is when nodes have community memberships, such that recovery is possible, as the SSC and SSS conditions determine it. Specifically, we show the following for a blockmodel graph (proof deferred in the Appendix).
Proposition 4.4.
If $G$ has two communities with $n_{1}$ and $n_{2}$ nodes respectively such that $n_{1}\geq n_{2}\geq 1$, and $X\in\{0,1\}^{n\times 2}$ is an embedding vector where each row $\bm{x}_{i}$ corresponds to a one-hot vector for the community of node $i$, then Algorithm 2 can recover $S$ perfectly as long as $|S|<\frac{n_{1}}{17}$.
For instance, when Proposition 4.4 is applied to the Polblogs dataset, it shows that $S$ can be fully recovered as long as $|S|\leq 9$. We can obtain a slightly worse bound and extend the result to a blockmodel graph with $K$ communities (proof in the Appendix).
Proposition 4.5.
If $G$ has $K\geq 2$ communities with sizes $n_{1}\geq n_{2}\geq\dots\geq n_{K}$ with $\left(\frac{16K}{16K+1}\right)\frac{n}{K}<n_{K}\leq\frac{n}{K}$ and $X\in\{0,1\}^{n\times K}$ is an embedding vector where $\bm{x}_{i}$ corresponds to an one-hot encoding of the community membership, then Algorithm 2 can recover $S$ perfectly as long as $|S|<\frac{1}{16K+1}n$.
In detail, Proposition 4.5 states that as long as the smallest community of the graph has size $\Theta(n/K)$ then the recovery of a set of size $|S|=O(n/K)$ is possible. If $\lvert S\rvert\gg n/K$ and $S$ contains all members of the smallest community, then robust regression can fail.
We show that in real-world datasets, Algorithm 2 can identify the set of deviators with high accuracy (cf. Figure 7). Specifically, in the real-world datasets we take $S$ to be a randomly sampled set of size $\lceil pn\rceil$ for $p\in\{0.05,0.1,0.15,0.2\}$ and the embeddings to be 128-dimenaional Node2Vec embeddings for Twitter and Reddit and community membership embeddings for the Pollblogs dataset. Algorithm 2 achieves low recovery error as well as high balanced accuracy score.
Figure 7. Reconstruction error and balanced accuracy for the robust regression problem presented in Algorithm 2. The $x$-axis shows the percentage of strategic agents. The left subfigure shows the recovery error, measured as $\frac{1}{n}\sum_{i\in[n]}\left|\frac{\hat{\bm{s}}_{i}-\bm{s}_{i}}{\bm{s}_{i}}\right|$, and the right subfigure measures the balanced accuracy between the recovered $\hat{S}$ and the true $S$. To construct the confidence intervals, for each size $|S|$ of the set $S$, we draw $S$ five times randomly from the vertex set $[n]$. For the Twitter and Reddit datasets, we have used 128-dimensional Node2vec embeddings. For the Polblogs dataset we have used the community membership (which corresponds to the political orientation) of each node, such that $\bm{x}_{i}=(1,0)$ corresponds to liberal and $\bm{x}_{i}=(0,1)$ corresponds to conservative, and (the true) $\bm{s}$ is such that $\bm{s}=X\bm{v}$ for $\bm{v}=(1,-1)^{T}$. We have also provided results using 128-dimensional spectral embeddings. We have set the recovery threshold for Torrent to be $|S|/n$, and the step size to be $\eta=1/\|\bar{X}\|_{2}^{2}$ where $\bar{X}$ is the min-max normalized embedding matrix.
5. Discussion and Conclusion
In this paper, we examine how opinions evolve in social networks, where individuals adjust their publicly stated views based on interactions with others and their inherent beliefs. In our model, strategically motivated users can distort these dynamics by misrepresenting their intrinsic opinions, often to advance conflicting objectives or promote rival narratives. We analyze the Nash Equilibrium of the resulting strategic interactions and empirically show — using diverse datasets from Twitter, Reddit, and Political Blogs — that such deceptive behavior intensifies polarization, fuels disagreement, and increases equilibrium costs. Additionally, we establish worst-case guarantees on the Price of Misreporting, akin to the Price of Anarchy, and introduce scalable learning algorithms to help platforms (i) detect opinion manipulation and (ii) identify the users responsible. Our algorithms perform effectively on real-world data, suggesting how platforms might mitigate the effects of strategic opinion shaping.
We conclude with some discussion of the implications of our work, and directions for future work.
Structural Platform Interventions.
We give algorithms for platforms to detect in strategic manipulation has occurred, and who is responsible (Section 4). Having done so, a platform might seek to mitigate the effects of strategic behavior. There are multiple plausible avenues to do so.
First, they may seek to reduce degree disparities in the network by algorithmically encouraging balanced connections, such as by suggesting users connect with those who have fewer connections, or reducing the visibility of central nodes (hubs). As we saw, both the upper bound on the $\mathsf{PoM}$ and the ratio $\|\bm{s}^{\prime}\|_{2}/\|\bm{s}\|_{2}$ depend on the largest eigenvalue $\lambda_{n}$ of the Laplacian, which scales with the maximum degree. This motivates interventions to balance the degree distribution.
Second, platforms can design strategy-proof mechanisms to incentivize the agents to report their true opinions. For resource allocation games on networks, it is known that the classical Vickrey–Clarke–Groves (VCG) mechanism is susceptible to adversarial behaviors such as collusion, motivating the need for different mechanisms (Chorppath et al., 2015). In the case of social networks, platforms have unique tools such as fines or banning of accounts to modify agents’ utility functions.
Future Work
First, it is not clear which measure of centrality should be used to identify the users who are most capable of manipulating others. As can be seen from Theorem 2.2, the influence of agent $i\in S$ should depend on the other members of $S$, as well as the spectral properties of the localized Laplacian matrices $L_{j}$ for $j\in S$ and susceptibility parameters $\alpha_{k}$ for $k\in[n]$.
Second, providing $\mathsf{PoM}$ bounds where $S$ can be any set of agents constitutes another interesting research direction, especially if platforms can assume that $S$ is a small fraction of all users. As we noted after Theorem 3.1, we believe that this would require restricted invertibility analysis of the matrices determining Nash equilibria.
Third, future work might consider different models of strategic manipulation. For example, one could consider a “feedback equilibrium” model (in the sense of dynamic games (Li et al., 2024)), in which agents $i\in S$ can report arbitrary $\bm{z}_{i}^{\prime}(t)$ at each timestep $t$, rather than following Eq. (1). This flexbility may give strategic agents more power to influence outcomes.
Appendix A Proofs
A.1. Proof of Proposition 4.4
Let $V_{1}$ correspond to the vertex set for community 1 and $V_{2}$ correspond to the vertex set for community 2. Let $I_{q}$ denote the $q\times q$ identity matrix, and $a_{i}=\lvert V_{i}\cap S\rvert$ for $i=1,2$. Then
$X_{S}^{T}X_{S}=\begin{pmatrix}I_{a_{1}}&0\\ 0&I_{a_{2}}\end{pmatrix}.$
Therefore $\lambda_{\max}(X_{S}^{T}X_{S})=\max\{|V_{1}\cap S|,|V_{2}\cap S|\}$ and $\lambda_{\min}(X_{S}^{T}X_{S})=\min\{|V_{1}\cap S|,|V_{1}\cap S|\}$.
First, we determine sufficient ranges of $\gamma$ and the value of $\Xi_{\gamma}$: Let $S$ be such that $|S|=\gamma n$. We have the following options:
  • $S\subseteq V_{1}$. Then $\lambda_{\max}(X_{S}^{T}X_{S})=\gamma n$.
  • $S\subseteq V_{2}$. Then $\lambda_{\max}(X_{S}^{T}X_{S})=\gamma n$.
  • $V_{1}\subseteq S$. Then $\lambda_{\max}(X_{S}^{T}X_{S})$ lies between $\gamma n/2$ and $\gamma n$.
  • $V_{2}\subseteq S$. Then $\lambda_{\max}(X_{S}^{T}X_{S})$ lies between $\gamma n/2$ and $\gamma n$.
  • If $S$ lies partially in $V_{1}$ and $V_{2}$, then $\lambda_{\max}(X_{S}^{T}X_{S})=\max\{(1-t)\gamma n,t\gamma n\}$ for some $t\in[0,1]$. Again, this is upper bounded by $\gamma n$.
The above yield $\Xi_{\gamma}=\gamma n$. To determine $\xi_{1-\gamma}$ we let $S$ be such that $|S|=(1-\gamma)n$. We have the following options:
  • $S\subseteq V_{1}$. Then $\lambda_{\min}(X_{S}^{T}X_{S})=0$.
  • $S\subseteq V_{2}$. Then $\lambda_{\min}(X_{S}^{T}X_{S})=0$.
  • $V_{1}\subseteq S$. Then $\lambda_{\min}(X_{S}^{T}X_{S})=\min\{n_{1},(1-\gamma)n-n_{1}\}=n_{1}\geq(1-\gamma)n-n_{2}$ since always $n_{1}\geq(1-\gamma)n/2$.
  • $V_{2}\subseteq S$. Then $\lambda_{\min}(X_{S}^{T}X_{S})=\min\{n_{2},(1-\gamma)n-n_{2}\}=(1-\gamma)n-n_{2}$ since $n_{2}\leq(1-\gamma)n/2$.
  • If $S$ lies partially in $V_{1}$ and $V_{2}$, then $\lambda_{\min}(X_{S}^{T}X_{S})=\min\{(1-t)(1-\gamma)n,t(1-\gamma)n\}$ for some $t\in[0,1]$. Again, this is lower bounded by $(1-\gamma)n-n_{2}$.
Therefore, for either $1-\gamma\leq n_{1}/n$ or $1-\gamma\leq n_{2}/n$ we have that $\xi_{1-\gamma}=(1-\gamma)n-n_{2}$. The final inequality corresponds to
$4\sqrt{\frac{\Xi_{\gamma}}{\xi_{1-\gamma}}}<1\iff\gamma<\frac{1}{17}-\frac{n_{2}}{17n}$
Combining the above we get two systems of inequalities. The first one corresponds to $1-\frac{n_{1}}{n}\leq\gamma<\frac{1}{17}-\frac{n_{2}}{17n}$ which holds for $n_{2}<1/18$ which is impossible since $n_{2}\geq 1$. The second one corresponds to $1-\frac{n_{2}}{n}\leq\gamma<\frac{1}{17}-\frac{n_{2}}{17n}$ which holds for $n_{2}>16/18$, which is always true.
Appendix B Proof of Proposition 4.5
First, we note that if $V_{1},\dots,V_{K}$ are the vertex sets and $S$ is a set of size $\gamma n$ the maximum eigenvalue equals to $\lambda_{\max}(X_{S}^{T}X_{S})=\max_{i\in[K]}|V_{i}\cap S|$ and is always at most $\gamma n$. So $\Xi_{\gamma}=\gamma n$. If $S$ is a set of size $(1-\gamma)n$, then the minimum eigenvalue $\lambda_{\min}(X_{S}^{T}X_{S})=\min_{i\in[K}|V_{i}\cap S|$ is maximized when $|V_{1}\cap S|=\dots=|V_{K}\cap S|=(1-\gamma)n/K$ which holds as long as $(1-\gamma)n/K\leq n_{K}$, so $\xi_{1-\gamma}=(1-\gamma)n/K$ as long as $\gamma\geq 1-n_{K}/nK$. Also the other condition is
$4\sqrt{\frac{\Xi_{\gamma}}{\xi_{1-\gamma}}}<1\iff\gamma<\frac{1}{16K+1}$
Finally, we must have $1/(16K+1)>1-Kn_{K}/K$ which yields $n_{K}>n/K(16K/(16K+1))$.
Appendix C Additional Figures
Figure 8. Distribution of centralities and degrees for the datasets
Figure 9. Running the experiments of Figure 3 for the Reddit dataset.
Figure 10. Running the experiments of Figure 3 for the Twitter dataset.